OpenStudy (anonymous):

sin(3theta) = -1 find the solution set. I only got the first radian right. which was (pie/2)

OpenStudy (dumbcow):

sin = -1 when angle is 3pi/2 plus multiples of 2pi 3pi/2 + 2pi = 7pi/2 --->theta = 7pi/6

OpenStudy (anonymous):

\[\text{lets see begin with breaking this up.}\\sin(3\theta)=sin(\theta +2\theta) \\sin(A+B)=sinAcosA+sinAcosB \\ proceed--->try \]

OpenStudy (anonymous):

and dumbcow how did you get that method of doing multiples of 2

OpenStudy (dumbcow):

i dont think mebs method applies to your problem

OpenStudy (dumbcow):

yes keep adding 2pi to get more solutions

OpenStudy (anonymous):

yea i noticed . it doesnt

OpenStudy (anonymous):

but where did you get 2pie from. ?

OpenStudy (dumbcow):

thats the num of radians in a circle

OpenStudy (anonymous):

hmmm... it is to tedious don't bother I was thinking of something else.

OpenStudy (anonymous):

its alright mebs. And dumbcow thanks I figured it out.